# Diffusion Beats Autoregressive in Data-Constrained Settings

### v1

@@ 4.2 Fitting Data-Constrained Scaling Laws
Figure 4 shows validation loss versus training tokens and reinforces the trend: AR models overfit with increased repetition, showing diverging loss curves. In contrast, diffusion models exhibit overlapping curves across repetitions, indicating no signs of overfitting and a very low decay rate with data reuse..
Figure 5 shows extrapolated training curves at large compute budgets for AR and diffusion models using our fitted scaling laws. For each setting, we use the compute-optimal model and dataset size derived from single-epoch scaling laws and extend training across multiple epochs. The dashed lines represent the ideal Chinchilla-style scaling behavior, where all training tokens are assumed to be unique. We find that for AR models, repeated data provides nearly the same benefit as fresh data only up to about 4 epochs. Beyond this point, additional repetition yields diminishing returns. In contrast, diffusion models continue to match the unique-data curve for up to 100 epochs, indicating a far greater capacity to benefit from repeated data in data-constrained regimes.
@@ 4.3 When to Use Diffusion over AR?
$\log_{10}(U)=0.460\cdot\log_{10}(C)-7.052,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{15}\cdot U^{2.174}.$
@@ 5 Discussion
We hypothesize that the key advantage stems from the use of random masking in diffusion models, which serves as a form of data augmentation. Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself. This broader distribution over prediction tasks encourages better generalization and more effective use of each training example. Analogous to image-based tasks where techniques like random cropping or color jittering boost generalization  caron2021emerging ; cubuk2020randaugment ; he2022masked , we believe the corruption process in diffusion plays a similar role. By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency.
@@ 8 Acknowledgement
We thank Alexander Li for his valuable insights and detailed feedback on the manuscript. We are also grateful to Zheyang Qin for his help in improving the figures. Finally, we thank Niklas Muennighoff, Simo Ryu and Colin Raffel for their helpful comments on the final draft of the paper. This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.

### v2

@@ 4.2 Fitting Data-Constrained Scaling Laws
Figure 4 shows validation loss versus training tokens using the compute budget of 1e19. The results reinforces the trend: AR models overfit with increased repetition, showing diverging loss curves. In contrast, diffusion models exhibit overlapping curves across repetitions, indicating no signs of overfitting and a very low decay rate with data reuse.
Figure 5 shows extrapolated training curves at large compute budgets. For each setting, we use the compute-optimal model and dataset size derived from single-epoch scaling laws for 1e19, 3e19 and 1e20. We then extend training to multiple epochs. The dashed lines represent the ideal Chinchilla-style scaling behavior, where all training tokens are assumed to be unique. We find that for AR models, repeated data provides nearly the same benefit as fresh data only up to about 4 epochs. Beyond this point, additional repetition yields diminishing returns. In contrast, diffusion models continue to match the unique-data curve for up to 100 epochs, indicating a far greater capacity to benefit from repeated data in data-constrained regimes.
@@ 4.3 When to Use Diffusion over AR?
$\log_{10}(U)=0.460\cdot\log_{10}(C)-7.052,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
@@ 5 Discussion
We hypothesize that the key advantage stems from the use of random factorization in diffusion models, which serves as a form of data augmentation. Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself. This broader distribution over prediction tasks encourages better generalization and more effective use of each training example. Analogous to image-based tasks where techniques like random cropping or color jittering boost generalization  caron2021emerging ; cubuk2020randaugment ; he2022masked , we believe the corruption process in diffusion plays a similar role. By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency.
@@ 8 Acknowledgement
We thank Alexander Li for his valuable insights and detailed feedback on the manuscript. We are also grateful to Zheyang Qin for his help in improving the figures. Finally, we thank Niklas Muennighoff, Simo Ryu and Colin Raffel for their helpful comments on the final draft of the paper. We thank Stella Biderman for pointing out the inconsistent unit used in the Critical Compute equation. This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.